Two objects (38.5 and 20.8 kg) are connected by a massless string that passes over a massless, frictionless pulley. The pulley hangs from the ceiling. Find the acceleration of the objects and the tension in the string acceleration tension m/s2

Could you please show your work as well? Thank you!
Answer #1

m2 ml m2g m1g

m1 > m2


m1 moves downwards and m2 moves upwards

for mass m1 = 38.5

Fnet = m1*a

T - m1*2g = -m1*a

T = m1*g - m1*a.............(1)


for mass m2 = 20.8

Fnet = m2*a

T - m2*g = m2*a


T = m2*g + m2*a .............(2)


from 1 & 2

m2*g + m2*a = m1*g - m1*a

m1*a + m2*a = m1*g - m2*g


acceleration a = (m1 - m2)*g/(m1+m2)


a = (38.5-20.8)*9.81/(38.5+20.8)


acceleration a = 2.93 m/s^2 <<<------------answer

===================


from equation 1

T = m1*g - m1*a


T = m1*(g-a)

T = 38.5*(9.81 - 2.93)

T = 265 N